OpenStudy (anonymous):

(cos^2x + sin^2x)/(cosx+sinx)^2 Please simplify thanks

OpenStudy (saifoo.khan):

\[\text{ you should just \end up with} \to \frac{1}{\sin^2x + 2\sin x \cos x + \cos ^2 x}\]

OpenStudy (sanchez9457):

you could also end up with: 1/1+2sinxcosx

OpenStudy (sanchez9457):

\[\frac{ 1 }{ 1+2sinxcosx }\]

OpenStudy (saifoo.khan):

Or also with: 1/(sin2x+1)

OpenStudy (saifoo.khan):

:D

OpenStudy (tkhunny):

OR \(\dfrac{1}{2\cdot\cos^{2}\left(x-\dfrac{\pi}{4}\right)}\)